OpenStudy (anonymous):

the line has slope 2 and passes through the points (-1, -6) and (2, -4)

OpenStudy (anonymous):

so whats the question? are you looking for the y intercept?

OpenStudy (anonymous):

im looking for an equation in slope intercept form

OpenStudy (y1066380):

use the point slope form (y1-y)=m(x1-x) m is the slope

OpenStudy (y1066380):

m=(-4+6)/(2+1) m=2/3 Put (2,-4) and m into (y1-y)=m(x1-x) (-4-y)=2/3(2-x) -12-3y=4-2x 2x-3y-16=0 is this correct?

OpenStudy (anonymous):

thanks